# Variational Adaptive Gaussian Decomposition: Scalable Quadrature-Free Time-Sliced Thawed Gaussian Dynamics

### 11 Mar 2026 v1

@@ III.1.2 Multidimensional Explorations
Pushing the semiclassical simulations to the full quantum levels of accuracy gets increasingly difficult with dimensionality. The time-slicing procedure, if done using quadrature, can become intractable very quickly. To study the scaling of VAGD with dimensions, consider the simple case of $d$ uncoupled Morse oscillators as an example:
@@
where all the centers at origin ($x_{0,i}=0.0$) and $\omega=$900\text{\,}{\mathrm{cm}}^{-1}$\$ and $\chi=0.005$ for all dimensions. While the scaling with dimensionality would of course be worse for cases where the different dimensions are correlated, this independent dimensions model allows us to show study the scaling in a relatively simplified setting. Our time-sliced TGA does not “know” that the dimensions can be separated and hence does not use that information explicitly. Though the neural networks can optimize to a certain extent for simplicity. We will explore correlated multidimensional models in terms of the double-well potentials.
@@
with zero initial momentum and ground-state width along each dimension. We start by demonstrating the deterioration of the quality of the standard TGA simulation for different dimensionality of the system in Fig. 5. A time-step of $\Delta t=$0.053\text{\,}\mathrm{fs}$\$ was used for all the calculations. Because the dimensions are independent and the displacement of the initial wave packet decreases with $d$, the degradation of quality of TGA is less pronounced for larger $d$.
Figure 6: Overlap of VAGD-TGA wave function with SOFT solution (solid, left $y$-axis) and number of VAGD trajectories (dashed, right $y$-axis) for multidimensional Morse models as a function of dimensionality.
Next, we perform the VAGD procedure every $25$ steps (segment length $\tau=25\Delta t$). At each decomposition boundary, the VAGD-reconstructed state is required to satisfy $F_{\text{thresh}}=0.9995$ with the propagated and the $W_{\text{min}}$ for individual Gaussian wave packets was set to 5.0. The results shown in Fig. 6 demonstrate the recovery of full quantum mechanical results. In all these cases, the number of trajectories saturated to the maximum limit allowed, $K$. However, there are a couple of things to be noted. First, the number of trajectories do not immediately converge to $K$. Depending upon the exact system setup, it can use less trajectories as well. This is the adaptive nature of the algorithm. The threshold $K$ that is required is also a function of the duration of simulation. The longer a wave packet propagates in an anharmonic potential, the more non-Gaussian character it develops, and the higher is the number of trajectories that are required.
Figure 7: Maximum allowed number of GWPs, $K$, used in the present multidimensional Morse simulations as a function of dimensionality.
Finally, in Fig. 7, we summarize the values of the maximum allowed number of GWPs, $K$, that were required to obtain the converged multidimensional Morse results up to a final time of $t_{\text{final}}=$212\text{\,}\mathrm{fs}$\$. For the present set of calculations, the observed values are $K=6$, $16$, $48$, and $96$ for $d=1$, 22, 33, and $4$, respectively.
It is important, however, to interpret Fig. 7 only as an empirical summary of the present numerical experiments. Therefore, we do not make any broader claim regarding a universal scaling law from these data alone. Nevertheless, the trend in Fig. 7 provides a useful indication of how the complexity of the adaptive Gaussian representation grows for these multidimensional Morse calculations.
@@ III.2 Double Well Potentials
As a next class of more complicated examples, we demonstrate VAGD method on the double-well potential. These systems are dynamically strongly correlated in time, in the sense that no single classical trajectory can capture the true dynamics. The deep tunneling observed in these systems cannot be truly described by simple single trajectory methods, and therefore, TGA also fails. Therefore time-sliced TGA has been shown to be able to be able to account for the tunneling event over [17]. First we show how the neural network approach is able to optimally represent the wave-function in one- and two- dimensions.
@@ III.2.1 One-Dimensional Problem
In addition to the overlap with SOFT results, following Ref. 17, we calculate the following correlation function,
@@
measuring the overlap of the time-propagated wave function and the mirror-image of the initial wave function. This is a measure of how well the approximation is able to capture the deep-tunneling events.
Figure 9: Number of trajectories for the 1D double well system.
@@
Notice in Fig. 8 (b), the TGA wave function does not tunnel. This is expected because of the purely classical nature of it. It is interesting that in our VAGD-TGA simulations, upto $K=8$, the wave function does not really tunnel. There is some fractional tunneling event. As we move from $K=8$ to $K=9$, there is a jump in the amount of tunneling which is captured. Qualitatively, all the essential features of $|\absolutevalue{C(t)}$ is achieved by a $K=9$ calculation, though the quantitative agreement is still missing. It is the second tunneling event back to the original well that is difficult. Using a maximum of $K=14$ trajectories, we start to get quantitative agreement of the correlation function upto a maximum time of 150. This is a significant improvement over the original TSTG method [17], which used 2048 trajectories in comparison to VAGD method using tens of trajectories.
In Fig. 9, we show the actual number of trajectories used as a function of time for different values of $K$. We see a difference from the previous examples. Here, the number of trajectories almost immediately jumps to the maximum value. This Hamiltonian is significantly more anharmonic than the Morse potentials considered earlier. Hence the necessity for a larger number of trajectories. Even the adaptive algorithm is unable to restrict the $N_{\text{traj}}(t)$ to very low values. If we had not capped it, the number would have been even larger. However, the perfect overlaps tell us that the simulations are converged with respect to $K$. Using larger values of $K$ would have resulted in the adaptive algorithm making the $N_{\text{traj}}(t)$ significantly larger than what is physically required.
@@ III.2.2 Tunneling in a Two-Dimensional Double-Well
Finally, we explore the 2D double well studied by Kong et al. [17]. The model potential couples two strongly correlated degrees of freedom and is given by
@@
As in the one-dimensional case, the quartic term generates a double-well structure along $x_{1}$, while the bilinear coupling term $(x_{1}x_{2}/2)$ induces a strong correlation between the coordinates. The initial state is chosen as a single normalized GWP localized in the left well along $x_{1}$ and centered at $\vec{q}_{0}=(-2\sqrt{\eta},0)$ with zero momentum along both directions,
@@
As expected, the TGA fails rapidly for this strongly correlated tunneling problem. Its overlap with the exact SOFT result drops quickly at early times and then exhibits large oscillations, indicating that a single thawed Gaussian is unable to represent the evolving correlated structure of the wave function. In contrast, both VAGD-TGA calculations retain substantially higher overlap throughout the propagation window. The $K=200$ result already provides a major improvement over TGA, while the $K=600$ calculation gives a further systematic gain and maintains better agreement at longer times and is nearly exact up to t=60a.u.t=60\,a.u. For the present setup, the benefit of allowing a larger Gaussian representation becomes especially clear beyond about $t\approx 80\,a.u.$, where the dynamics develops stronger tunneling and interference character.
@@
To examine this more directly, in Fig. 11 we show the population that has tunneled to the open right half-plane,
@@ IV Conclusion
VAGD provides a variationally optimal decomposition of a given wave function in terms of GWPs. Its main advantage is that the effective “basis” is optimized specifically for the wave function under consideration, through an autoencoder–decoder architecture. In combination with time-sliced TGA (VAGD–TGA), the method has been tested on the Morse potential and on one- and two-dimensional tunneling problems. Dimensional scaling was further examined using independent Morse oscillators up to four dimensions, where the computational cost was found to grow only mildly with dimensionality. Our results suggest a polynomial scaling behavior, although the precise performance will naturally depend on the system under study. Nevertheless, VAGD promises a substantial improvement over the exponential scaling typically encountered in conventional approaches.
@@
Beyond its practical efficiency, VAGD additionally offers an appealing conceptual perspective on semiclassical dynamics. The number of Gaussian wave packets required in the decomposition naturally reflects the degree of non-classical structure present in the evolving wave function. Systems exhibiting largely classical behavior can be represented with only a small number of trajectories, whereas strongly quantum phenomena, such as interference and tunneling, require a richer Gaussian representation. In this sense, the adaptive Gaussian basis produced by VAGD provides a direct measure of the intrinsic quantum complexity of the dynamics. Looking forward, the approach opens the door to scalable semiclassical simulations in higher-dimensional systems and even ab initio molecular models, where the combination of time-sliced dynamics with adaptive Gaussian decompositions may provide an efficient route to near–quantum-accurate simulations.

### 12 Mar 2026 v2

@@ III.1.2 Multidimensional Explorations
Pushing the semiclassical simulations to the full quantum levels of accuracy gets increasingly difficult with dimensionality. The time-slicing procedure, if done using quadrature, can very quickly become intractable. To study the scaling of VAGD with dimensions, consider the simple case of $d$ uncoupled Morse oscillators as an example:
@@
where all the centers at origin ($x_{0,i}=0.0$) and $\omega=$900\text{\,}{\mathrm{cm}}^{-1}$\$ and $\chi=0.005$ for all dimensions. While the scaling with dimensionality would of course be worse for cases where the different dimensions are correlated, this independent dimensional model allows us to study the scaling in a relatively simplified setting. Our time-sliced TGA does not “know” that the dimensions can be separated and hence does not use that information explicitly. We will explore correlated multidimensional models in terms of the double-well potentials.
@@
with zero initial momentum. We start by demonstrating the deterioration of the quality of the standard TGA simulation for different dimensionality of the system in Fig. 5. A time-step of $\Delta t=$0.053\text{\,}\mathrm{fs}$\$ was used for all the calculations. Because the dimensions are independent and the displacement of the initial wave packet along each dimension decreases with $d$, the degradation of quality of TGA is less pronounced for larger $d$.
Figure 6: Overlap of VAGD-TGA wave function with SOFT solution (solid, left $y$-axis) and number of trajectories (dashed, right $y$-axis) for multidimensional Morse models as a function of dimensionality.
Next, we perform the VAGD procedure every 25 steps. At each decomposition boundary, the VAGD-reconstructed state is required to satisfy $F_{\text{thresh}}=0.9995$ with the propagated and the $W_{\text{min}}$ for individual Gaussian wave packets was set to 5.0 along each dimension. The results shown in Fig. 6 demonstrate the recovery of full quantum mechanical results. In all these cases, the number of trajectories saturated to the maximum limit allowed, $K$. However, there are a couple of things to be noted. First, the number of trajectories do not immediately converge to $K$. Depending upon the exact system setup, it can use less trajectories as well. This is the adaptive nature of the algorithm. The threshold $K$ that is required is also a function of the duration of simulation. The longer a wave packet propagates in an anharmonic potential, the more non-Gaussian character it develops, and the higher is the number of trajectories that are required.
Figure 7: Number of trajectories required for converged results at $t=$212\text{\,}\mathrm{fs}$\$ for different numbers of dimensions.
Finally, in Fig. 7, we summarize the number of trajectories required to obtain the converged results at a final time of $t_{\text{final}}=$212\text{\,}\mathrm{fs}$\$. It is important, however, to interpret Fig. 7 only as an empirical summary of the present numerical experiments. Therefore, we do not make any broader claim regarding a universal scaling law from these data alone. Nevertheless, the trend provides a useful indication of how the complexity of the adaptive Gaussian representation grows for these multidimensional Morse calculations.
@@ III.2 Double Well Potentials
As a next class of more complicated examples, we demonstrate VAGD method on the double-well potential. These systems are dynamically strongly correlated in time, in the sense that no single classical trajectory can capture the true dynamics. The deep tunneling observed in these systems cannot be described by simple single trajectory methods, and therefore, TGA fails. Time-sliced TGA has been shown to be able to account for the tunneling event [17]. We demonstrate how VAGD is able to optimally represent the wave-function in one- and two- dimensions.
@@ III.2.1 One-Dimensional Problem
Figure 9: Number of trajectories for the 1D double well system.
In addition to the overlap with SOFT results, following Ref. 17, we correlate the mirror-image of the initial wave function with the time-propagated one,
@@
measuring the extent of tunneling captured by the various levels of approximation.
@@
Notice in Fig. 8 (b), the TGA wave function does not tunnel at all. This is expected because of its purely classical nature. It is interesting that in our VAGD-TGA simulations, upto $K=8$, the wave function does not really tunnel. As we move from $K=8$ to $K=9$, there is a jump in the amount of tunneling captured. Qualitatively, all the essential features of $|\absolutevalue{C(t)}$ is achieved by a $K=9$ calculation, though the quantitative agreement is still missing. It is the second tunneling event back to the original well that proves to be even more challenging. Using a maximum of $K=14$ trajectories, we start to get quantitative agreement of the correlation function upto a maximum time of $t=150$. This is a significant improvement over the original TSTG method [17], which used 2048 trajectories in comparison to VAGD method using tens of trajectories.
In Fig. 9, we show the actual number of trajectories used as a function of time for different values of $K$. We see a difference from the previous examples. Here, the number of trajectories almost immediately jumps to the maximum value owing to the significantly greater degree of anharmonicity. Even the adaptive algorithm is unable to effectively restrict the $N_{\text{traj}}(t)$ to extremely low values. If we had not capped it, the number would have been even larger. However, the perfect overlaps tell us that the simulations are converged with respect to $K$. Using larger values of $K$ would have resulted in the adaptive algorithm making the $N_{\text{traj}}(t)$ significantly higher than what is required.
@@ III.2.2 Tunneling in a Two-Dimensional Double-Well
Finally, we explore the 2D double well studied by Kong et al. [17] The model potential couples two strongly correlated degrees of freedom and is given by
@@
As in the one-dimensional case, the quartic term generates a double-well structure along $x_{1}$, while the bilinear coupling term $(x_{1}x_{2}/2)$ induces a strong correlation between the coordinates. The initial state is chosen as a single normalized GWP localized in the left well along $x_{1}$,
@@
As expected, TGA fails rapidly for this strongly correlated tunneling problem. Its overlap with the exact SOFT result drops quickly at early times and then exhibits large oscillations, indicating that a single thawed Gaussian is unable to represent the evolving correlated structure of the wave function. In contrast, both VAGD-TGA calculations retain substantially higher overlap throughout the propagation window. The $K=200$ result already provides a major improvement over TGA, while the $K=600$ calculation gives a further systematic gain and maintains better agreement at longer times and is exact up to t=60a.u.t=60\,a.u. For the present setup, the benefit of allowing a larger Gaussian representation becomes especially clear beyond about $t\approx 80\,a.u.$, where the dynamics develops stronger tunneling and interference character.
@@
To examine this more directly, in Fig. 11 we show the population that has tunneled to the right half-plane,
@@ IV Conclusion
VAGD provides a variationally optimal decomposition of a given wave function in terms of GWPs. Its main advantage is that the effective “basis” is optimized specifically for the wave function under consideration through an autoencoder–decoder architecture. In combination with time-sliced TGA (VAGD–TGA), the method has been tested on the Morse potential and on one- and two-dimensional tunneling problems. Dimensional scaling was further examined using independent Morse oscillators up to four dimensions, where the computational cost was found to grow only mildly with dimensionality. Our results suggest a polynomial scaling behavior, although the precise performance can not yet be commented on from the numerical explorations. Nevertheless, VAGD promises a substantial improvement over the exponential scaling typically encountered in conventional approaches.
@@
Beyond its practical efficiency, VAGD additionally offers an appealing conceptual perspective on semiclassical dynamics. The number of Gaussian wave packets required in the decomposition naturally reflects the degree of non-classical structure present in the evolving wave function. Systems exhibiting largely classical behavior can be represented with only a small number of trajectories, whereas strongly quantum phenomena, such as interference and tunneling, require richer Gaussian representations. In this sense, the adaptive Gaussian basis produced by VAGD provides a direct measure of the intrinsic quantum complexity of the dynamics. Looking forward, the approach opens the door to scalable semiclassical simulations in higher-dimensional systems and even ab initio molecular models, where the combination of time-sliced dynamics with adaptive Gaussian decompositions may provide an efficient route to near–quantum-accurate simulations.